OpenStudy (anonymous):

Find an equation of the circle that satisfies the stated conditions. (Give your answer in standard notation.) Tangent to both axes, center in the second quadrant, radius 3

OpenStudy (anonymous):

*

OpenStudy (anonymous):

Equation of a circle \[ (x+a)^2 + (y+b)^2 = r^2\] We make r = 3 so that it can have radius 3 \[ (x+a)^2 + (y+b)^2 = 3^2\] Then we translate it up and left by the radius (3) to make it tangent to each axis \[ (x+3)^2 + (y-3)^2 = 3^2\] Observe: http://www.wolframalpha.com/input/?i=%28x%2B3%29%5E2+%2B+%28y-3%29%5E2+%3D+3%5E2

OpenStudy (anonymous):

Remember that increasing a shifts it left and increasing b shifts it down.

OpenStudy (anonymous):

thanks :)